OpenStudy (anonymous):

the difference between seven times one number and three times a second number is 25. The sum of twice the first and five times the second is 95, Find the numbers

OpenStudy (anonymous):

7x-3y=25 2x+5y=95 System of equations, do you know how to solve?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

You want to solve for one variable in terms of the other.

OpenStudy (anonymous):

7x=3y+25 \[x=\frac{3}{7}y+\frac{25}{7}\]

OpenStudy (anonymous):

You then plug this into the other equation

OpenStudy (anonymous):

\[2(\frac{3}{7}y+\frac{25}{7})+5y=95\]

OpenStudy (anonymous):

\[\frac{6}{7}y+\frac{50}{7}+5y=95\]

OpenStudy (anonymous):

Multiply by 7

OpenStudy (anonymous):

\[6y+50+35y=665\]

OpenStudy (anonymous):

\[41y=615, y=15\]

OpenStudy (anonymous):

7x-3(15)=95 7x=140 x=20

OpenStudy (anonymous):

yur so kind thank you

OpenStudy (anonymous):

You're welcome